# Dwarf Galaxies in the MATLAS Survey: Hubble Space Telescope Observations of Nuclear Star Clusters

### v1

@@ Appendix B NSCs color profiles
Appendix C GCs properties
Table 2 presents the structural and photometric properties of the 142 marginally resolved GCs identified in 17 dwarf galaxies within a 25 Mpc distance (Marleau et al. 2024). For each GC, the table includes its RA/Dec coordinates, radial distance () in terms of galaxy Re,gal, absolute magnitude (MV,GC), color, effective radius (Re,GC), stellar mass (M), as well as the host galaxy with its distance. The last column indicates the band from which the parameters were derived. If the parameters are based on one band, 2Re,GC aperture photometry was applied in both bands to derive the color. When both bands were used, Re,GC corresponds to the average value of both bands. The GCs are sorted by increasing RA.

### v2

@@ Appendix B NSC color profiles
Appendix C GC properties
Table 2 presents the structural and photometric properties of the 142 marginally resolved GCs identified in 17 dwarf galaxies within a 25 Mpc distance (Marleau2024). For each GC, the table includes its RA/Dec coordinates, radial distance () in terms of galaxy Re,gal, absolute magnitude (MV,GC), color, effective radius (Re,GC), stellar mass (M), as well as the host galaxy with its distance. The last column indicates the band from which the parameters were derived. If the parameters are based on one band, 2Re,GC aperture photometry was applied in both bands to derive the color. When both bands were used, Re,GC corresponds to the average value of both bands. The GCs are sorted by increasing RA.
@@ Appendix C GC properties
∗ GC Re smaller than one pixel.
Distance measurement from aheesters2023, bSDSS DR13 database (Albareti2017; Blanton2017), and cMueller2021.
@@
∗ GC Re smaller than one pixel.
Distance measurement from aheesters2023, bSDSS DR13 database (Albareti2017; Blanton2017), and cMueller2021.